# Each of those eight coalition logics is also determined by six other kinds of models

### v2

@@ Abstract
Coalition Logic is a central logic in logical research on strategic reasoning. In two recent papers, Li and Ju argued that generally, concurrent game models, models of Coalition Logic, have three too strong assumptions: seriality, independence of agents, and determinism. They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions are met. In this paper, we show that each of the eight logics is also determined by six other kinds of models, with respective properties, that is, single-coalition-first action models, single-coalition-first neighborhood models, clear grand-coalition-first action models, clear single-coalition-first neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first neighborhood models.
@@ 1.1 Eight coalition logics determined by eight classes of general concurrent game models
Coalition Logic $\mathsf{CL}$, proposed by Pauly ([10, 11]), is a central logic in logical research on strategic reasoning.
The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle[\mathrm{C}]\rangle\phi$, indicating some available joint action of the coalition $\mathrm{C}$ ensures $\phi$.
$\mathsf{CL}$ has two kinds of models.
The first kind is concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents, who can form coalitions; for every coalition, there is an availability function, specifying available joint actions of the coalition at states; for every coalition, there is an outcome function, specifying possible outcome states of joint actions of the coalition. The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
The second kind is alpha neighborhood models. Alpha neighborhood models do not have actions. Consequently, they do not have availability functions and outcome functions. Instead, for every coalition, there is a neighborhood function that specifies the alpha powers of the coalition. An alpha power is a set of states into which the coalition can force the next moment to fall. The formula $\langle[\mathrm{C}]\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
[9] argued that concurrent game models have three too strong assumptions: seriality, the independence of agents, and determinism. Based on general concurrent game models, which do not have the three assumptions, they proposed a Minimal Coalition Logic $\mathsf{MCL}$. Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some kinds of situations. Considering which of the three properties we want to keep, there are eight coalition logics in total. [8] showed the completeness of the eight logics in a uniform way.
@@ 1.2 Our work
In this paper, we call general concurrent game models grand-coalition-first action models. In this work, we show that each of the eight coalition logics is also determined by six other kinds of models, with respective properties.
@@
The first kind is single-coalition-first action models. In this kind of models, the outcome functions for single coalitions determine the outcome functions for other coalitions. This kind of models makes good sense for situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other. The second kind is single-coalition-first neighborhood models. In this kind of models, the neighborhood function for a coalition specifies the actual powers of the coalition. An actual power is a minimal scope into which the coalition can force the next moment to fall. In this kind of models, the neighborhood functions for single coalitions determine the neighborhood functions for other coalitions.
@@
The third kind is clear grand-coalition-first action models. In this kind of models, different joint actions of the grand coalition have different outcomes. Clear grand-coalition-first action models are quite common in game theory. The fourth kind is clear single-coalition-first neighborhood models. In this kind of models, the actual powers of an agent are pairwise disjoint.
@@
The fifth kind is tree-like grand-coalition-first action models. In this kind of models, every state has a determined history. The sixth kind is tree-like single-coalition-first neighborhood models. In this kind of models, every state has a determined history, too.
The rest of the paper is structured as follows. In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from action semantics to neighborhood semantics. In Section 3, we introduce the eight coalition logics determined by eight classes of grand-coalition-first action models. In Section 4, we define single-coalition-first action models and show that they are grand-coalition-first action models; we define single-coalition-first neighborhood models and show that they can represent single-coalition-first action models. In Section 5, we define clear grand-coalition-first action models and show they are single-coalition-first action models; we define clear single-coalition-first neighborhood models and show that they can represent clear grand-coalition-first action models. In Section 6, we define tree-like grand-coalition-first action models and show that they are clear grand-coalition-first action models; we define tree-like single-coalition-first neighborhood models and show that they can represent tree-like grand-coalition-first action models. In Section 7, we show that each of the eight coalition logics is determined by the six kinds of models with respective properties. We conclude in Section 8.
@@ 2.1 Language
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle[\mathrm{C}]\rangle\phi$
The formula $\langle[\mathrm{C}]\rangle\phi$ indicates some available joint action of $\mathrm{C}$ ensures $\phi$. The propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ are defined as usual.
@@ 2.2 Action semantics
An action model is a tuple $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
@@
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an availability function for $\mathrm{C}$.
@@
Different coalition logics might put different constraints on action models.
@@
$\mathtt{AM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{AM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{AM},s\Vdash\phi$
$\mathtt{AM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{AM},s\Vdash\phi$ and $\mathtt{AM},s\Vdash\psi$
$\mathtt{AM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\phi$
@@ 2.3 Neighborhood semantics
Definition 4 (Neighborhood models and alpha neighborhood models).
A neighborhood model is a tuple $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
@@
A neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is an alpha neighborhood model if for every $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is closed under supersets.
Different coalition logics might put different constraints on neighborhood models.
@@
$\mathtt{NM},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{NM},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{NM},s\Vdash\phi$
$\mathtt{NM},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{NM},s\Vdash\phi$ and $\mathtt{NM},s\Vdash\psi$
$\mathtt{NM},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\phi$
The following fact, which is easy to verify, indicates that we can transform a pointed neighborhood model to a pointed alpha neighborhood model without changing the set of formulas it satisfies.
Fact 1.
Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be a neighborhood model and $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an alpha neighborhood model such that for all $\mathrm{C}\subseteq\mathtt{AG}$ and $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is the closure of $\mathtt{nei}_{\mathrm{C}}(s)$ under supersets.
Then, for every $s\in\mathtt{ST}$ and $\phi\in\Phi$, $\mathtt{NM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
@@ 2.4 Transformation of action semantics to neighborhood semantics
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
@@
Actual effectivity functions are more transparent in the following sense: each element of $\mathtt{AE}_{\mathrm{C}}(s)$ can be understood as a class of equivalent actions.
Definition 7 (Action models $z$-representable by neighborhood models and $\alpha$-representable by alpha neighborhood models).
An action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is $z$-representable by a neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
A class of action models $\mathbf{AM}$ is $z$-representable by a class of neighborhood models $\mathbf{NM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by a neighborhood mode in $\mathbf{NM}$, and (2) every neighborhood model in $\mathbf{NM}$ $z$-represents an action mode in $\mathbf{AM}$.
An action model $\mathtt{AM}$ is $\alpha$-representable by an alpha neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{LE}_{\mathrm{C}}=\mathtt{nei}^{\alpha}_{\mathrm{C}}$.
A class of action models $\mathbf{AM}$ is $\alpha$-representable by a class of alpha neighborhood models $\mathbf{NM}^{\alpha}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}^{\alpha}$, and (2) every alpha neighborhood model in $\mathbf{NM}^{\alpha}$ $\alpha$-represents an action mode in $\mathbf{AM}$.
Intuitively, that a neighborhood model $\mathtt{NM}$ represents an action model $\mathtt{AM}$ means that $\mathtt{NM}$ contains all the information of $\mathtt{AM}$ to evaluate formulas of $\Phi$.
@@
    For every action model $\mathtt{AM}$ and neighborhood model $\mathtt{NM}$, if $\mathtt{AM}$ is $z$-representable by $\mathtt{NM}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
@@
    For every action model $\mathtt{AM}$ and alpha neighborhood model $\mathtt{NM}^{\alpha}$, if $\mathtt{AM}$ is $\alpha$-representable by $\mathtt{NM}^{\alpha}$, then for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
    Consequently, for every class of action models $\mathbf{AM}$ and class of alpha neighborhood model $\mathbf{NM}^{\alpha}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}^{\alpha}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}^{\alpha}$.
3 Eight coalition logics determined by eight classes of grand-coalition-first action models

### v3

# Seven kinds of equivalent models

@@ Abstract
Coalition Logic is an important logic in logical research on strategic reasoning. In two recent papers, Li and Ju argued that generally, concurrent game models, models of Coalition Logic, have three too strong assumptions: seriality, independence of agents, and determinism. They presented eight coalition logics based on eight classes of general concurrent game models, determined by which of the three assumptions are met. In this paper, we show that each of the eight sets of valid formulas of the eight logics is determined by six other kinds of models, that is, single-coalition-first action models, single-coalition-first actual neighborhood models, clear grand-coalition-first action models, clear single-coalition-first actual neighborhood models, tree-like grand-coalition-first action models, and tree-like single-coalition-first actual neighborhood models.
@@ 1 Introduction
1.1 Coalition Logic and its two kinds of semantics
Coalition Logic $\mathsf{CL}$, proposed by Pauly [pauly_modal_2002], is an important logic in logical research on strategic reasoning. Many logics in this field are extensions of $\mathsf{CL}$. For example, Alternating-time Temporal Logic $\mathsf{ATL}$ [alur_alternating-time_2002] is a temporal extension of $\mathsf{CL}$, and Strategy Logic $\mathsf{SL}$ [mogavero_reasoning_2014] is an extension of $\mathsf{CL}$, whose language has quantifiers for and names of strategies. We refer to [benthem_models_2015], [agotnes_knowledge_2015] and [sep-logic-power-games] for overviews of the area.
The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$, indicating some available joint action of coalition $\mathrm{C}$ ensures $\phi$. $\mathsf{CL}$ has two kinds of equivalent semantics [pauly_modal_2002, goranko_strategic_2013]: action semantics and neighborhood semantics.
In action semantics, models are concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents, who can form coalitions; for every coalition, there is an availability function, specifying available joint actions of the coalition at states; for every coalition, there is an outcome function, specifying possible outcome states of joint actions of the coalition. The formula $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{C}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
In neighborhood semantics, models are alpha neighborhood models. Alpha neighborhood models do not have actions. Consequently, they do not have availability functions and outcome functions. Instead, for every coalition, there is a neighborhood function that specifies the alpha powers of the coalition. An alpha power is a set of states into which the coalition can force the next moment to fall [moulin_cores_1982]. The formula $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ is true at a state in an alpha neighborhood model if $\mathrm{C}$ has an alpha power such that $\phi$ is true at every state of the power.
Concurrent game models and alpha neighborhood models of $\mathsf{CL}$ have their respective advantages and disadvantages. Concurrent game models are more intuitive. However, they contain information for which the language of $\mathsf{CL}$ is blind, that is, actions. Neighborhood models are simpler: they only keep the needed information for evaluating formulas of $\mathsf{CL}$. Some work in the line of $\mathsf{CL}$, such as [agotnes_coalition_2018], uses neighborhood models
@@ 1.2 Eight coalition logics determined by eight classes of general concurrent game models
Concurrent game models have three assumptions, among others. The first one is seriality: every coalition always has an available joint action. The second one is the independence of agents: the merge of two available joint actions of two disjoint coalitions is always an available joint action of the union of the two coalitions. The third one is determinism: every available joint action of the grand coalition has a unique outcome. The three assumptions are also reflected in the alpha neighborhood models of $\mathsf{CL}$.
In the agency literature, some authors have used models in which one or more of the three assumptions do not hold. In the tradition of coalition logic, Jiang and Naumov [JIANG2022103727] used non-deterministic concurrent game models. In the tradition of STIT (see to it that) logic, Sergot [sergot_examples_2014, sergot_actual_2022] argued that neither the independence of agents nor determinism commonly holds. At the intersection of $\mathsf{ATL}$ and STIT logic, Boudou and Lorini [10.5555/3237383.3237443] utilized non-deterministic concurrent game models. In the tradition of Propositional Dynamic Logic, Royakkers and Hughes [royakkers_blame_2020] argued that the independence of agents does not hold generally and used models in which the independence of agents and determinism fail.
Against the agency literature, Li and Ju [li_minimal_2025] argued that the three assumptions are too strong. Here, we briefly mention their arguments. Usually, games terminate when reaching specific states. For example, a paper-scissors-rock game terminates when a winner is decided. Intuitively, in terminating states, players have no available actions. Thus, seriality might not hold. There are many situations in which whether a coalition can perform an action is conditional on other agents’ actions at the same time. Here is an example. There are two agents in a room, $a$ and $b$, and only one chair. Agent $a$ can sit, and agent $b$ can sit, but they cannot sit at the same time. The independence of agents fails in these situations. In many situations, some joint actions of all behaving agents have more than one outcome state. The following example, from [sergot_examples_2014], can illustrate this. A vase stands on a table. There is an agent $a$ who can lift or lower the end of the table. If the table tilts, the vase might fall, and if it falls, it might break. Therefore, determinism might not hold.
Based on general concurrent game models, which do not have the three assumptions, Li and Ju [li_minimal_2025] proposed a Minimal Coalition Logic $\mathsf{MCL}$. Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in some situations. Considering which of the three properties we want to keep, there are eight coalition logics in total. Li and Ju [li_completeness_2024] showed the completeness of the eight logics in a uniform way.
@@ 1.3 Our work
The eight coalition logics have eight sets of valid formulas. In this work, we show the following: for each of the eight sets of valid formulas, (1) it is also determined by three other classes of action models, and (2) it is determined by three other classes of actual neighborhood models. These different kinds of models make good sense for representing coalitional powers. Our work makes it clear what the logics with respect to these kinds of models are.
Action models are those like concurrent game models. In actual neighborhood models, for every coalition, there is a neighborhood function that specifies actual powers of the coalition. An actual power is a set of states such that the coalition has a choice such that (1) the choice forces the next moment to fall into the set, and (2) every state in the set is compatible with the choice [benthem_new_2019].
To ease understanding the differences among the six kinds of models, in this paper, we call general concurrent game models grand-coalition-first action models. The reason is that in general concurrent game models, the outcome functions of all coalitions can be derived from the outcome functions of the grand coalition.
The six kinds of models are as follows:
  1. 1.
    1. (a)
@@
      The first kind is single-coalition-first action models. In this kind of models, the outcome functions for single coalitions determine those for other coalitions. This kind of models makes good sense for situations where every agent can change a component of the state, and different components of the state that can be changed by different agents are independent of each other.
    2. (b)
      The second kind is single-coalition-first actual neighborhood models. In this kind of models, the neighborhood functions for single coalitions determine those for other coalitions.
  2. 2.
    1. (a)
@@
      The third kind is clear grand-coalition-first action models. In this kind of models, different joint actions of the grand coalition have different outcomes. Clear grand-coalition-first action models are quite common in game theory.
    2. (b)
      The fourth kind is clear single-coalition-first actual neighborhood models. In this kind of models, the actual powers of an agent are pairwise disjoint.
  3. 3.
    1. (a)
@@
      The fifth kind is tree-like grand-coalition-first action models. In this kind of models, every state has a determined history.
    2. (b)
      The sixth kind is tree-like single-coalition-first actual neighborhood models. In this kind of models, every state has a determined history, too.
This work offers new models for logics like Alternating-time Temporal Logic or STIT logic, which will be our future work.
Structure of the paper
  • In Section 2, we present some general settings of coalition logics, including their language, their action semantics and neighborhood semantics, and the transformation from action semantics to neighborhood semantics.
  • In Section 3, we introduce the eight coalition logics determined by eight classes of grand-coalition-first action models.
  • In Section 4, we define single-coalition-first action models and show that they are grand-coalition-first action models.
    We define single-coalition-first actual neighborhood models and show that they can represent single-coalition-first action models.
  • In Section 5, we define clear grand-coalition-first action models and show that they are single-coalition-first action models.
    We define clear single-coalition-first actual neighborhood models and show that they can represent clear grand-coalition-first action models.
  • In Section 6, we define tree-like grand-coalition-first action models and show that they are clear grand-coalition-first action models.
    We define tree-like single-coalition-first actual neighborhood models and show that they can represent tree-like grand-coalition-first action models.
  • In Section 7, we show that each of the eight coalition logics is determined by the six kinds of models with respective properties.
  • We conclude in Section 8.
@@ 2.1 Language
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$
The formula $\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ indicates some available joint action of $\mathrm{C}$ ensures $\phi$. Here, the notation of modalities differs from their notation in the literature. We do this to indicate the iteration of quantifiers in their meaning. The propositional connectives $\bot,\lor,\rightarrow$ and $\leftrightarrow$ are defined as usual.
@@ 2.2 Action semantics
An action model is a tuple $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
@@
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{av}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{JA}_{\mathrm{C}})$ is an availability function for $\mathrm{C}$ meeting the following condition: for all $s\in\mathtt{ST}$, $\mathtt{av}_{\mathrm{C}}(s)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})\neq\emptyset\}$.
@@
Different coalition logics might put various constraints on action models.
@@
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model. Then truth conditions of formulas in $\Phi$ at states $s$ are defined recursively as follows:
@@
$\mathtt{AM},s\Vdash p$ $\Leftrightarrow$ $p\in\mathsf{label}(s)$
$\mathtt{AM},s\Vdash\neg\phi$ $\Leftrightarrow$ not $\mathtt{AM},s\Vdash\phi$
$\mathtt{AM},s\Vdash\phi\land\psi$ $\Leftrightarrow$ $\mathtt{AM},s\Vdash\phi$ and $\mathtt{AM},s\Vdash\psi$
$\mathtt{AM},s\Vdash\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ $\Leftrightarrow$ there is $\sigma_{{\mathrm{C}}}\in\mathtt{av}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{{\mathrm{C}}})$, $\mathtt{AM},t\Vdash\phi$
@@ 2.3 Neighborhood semantics
Definition 4 (Actual neighborhood models).
An actual neighborhood model is a tuple $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
@@
Intuitively, $\mathtt{nei}_{\mathrm{C}}(s)$ consists of actual powers of $\mathrm{C}$ at $s$.
Definition 5 (Alpha neighborhood models).
An alpha neighborhood model is a tuple $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$, where:
  • $\mathtt{ST}$ is a nonempty set of states.
  • for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}:\mathtt{ST}\rightarrow\mathcal{P}(\mathcal{P}(\mathtt{ST}))$ is a neighborhood function for $\mathrm{C}$ such that for every $s\in\mathtt{ST}$, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ is closed under supersets.
Intuitively, $\mathtt{nei}^{\alpha}_{\mathrm{C}}(s)$ consists of alpha powers of $\mathrm{C}$ at $s$.
Different coalition logics might put various constraints on neighborhood models.
@@
Let $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an actual/alpha neighborhood model. Then truth conditions of formulas in $\Phi$ at states $s$ are defined recursively as follows:
@@
$\mathtt{NM},s\Vdash p$ $\Leftrightarrow$ $p\in\mathsf{label}(s)$
$\mathtt{NM},s\Vdash\neg\phi$ $\Leftrightarrow$ not $\mathtt{NM},s\Vdash\phi$
$\mathtt{NM},s\Vdash\phi\land\psi$ $\Leftrightarrow$ $\mathtt{NM},s\Vdash\phi$ and $\mathtt{NM},s\Vdash\psi$
$\mathtt{NM},s\Vdash\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\phi$ $\Leftrightarrow$ there is $\mathrm{Y}\in\mathtt{nei}_{\mathrm{C}}(s)$ such that for all $t\in\mathrm{Y}$, $\mathtt{NM},t\Vdash\phi$
@@ 2.4 Transformation of action semantics to neighborhood semantics
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
@@
Actual effectivity functions are more transparent than alpha effectivity functions in the following sense: every element $\mathrm{Y}$ of $\mathtt{AE}_{\mathrm{C}}(s)$ can be understood as a class of equivalent actions such that $\mathrm{Y}$ is the set of their possible outcome states at $s$.
The following fact, which is easy to verify, indicates that we can transform an action model to an actual/alpha neighborhood model without changing the satisfaction of formulas:
Fact 1.
Let $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ be an action model.
  • Then $\mathtt{NM}=(\mathtt{ST},\{\mathtt{AE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is an actual neighborhood model such that for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM},s\Vdash\phi$.
  • Then $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{LE}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is an alpha neighborhood model such that for every state $s$ of $\mathtt{AM}$ and formula $\phi$ in $\Phi$, $\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{NM}^{\alpha},s\Vdash\phi$.
Definition 8 (Action models $z$-representable by actual neighborhood models and $\alpha$-representable by alpha neighborhood models).
  • An action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is $z$-representable by an actual neighborhood model $\mathtt{NM}=(\mathtt{ST},\{\mathtt{nei}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{AE}_{\mathrm{C}}=\mathtt{nei}_{\mathrm{C}}$.
    A class of action models $\mathbf{AM}$ is $z$-representable by a class of actual neighborhood models $\mathbf{NM}$ if (1) every action model in $\mathbf{AM}$ is $z$-representable by an actual neighborhood model in $\mathbf{NM}$, and (2) every actual neighborhood model in $\mathbf{NM}$ $z$-represents an action model in $\mathbf{AM}$.
  • An action model $\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ is $\alpha$-representable by an alpha neighborhood model $\mathtt{NM}^{\alpha}=(\mathtt{ST},\{\mathtt{nei}^{\alpha}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ if for every $\mathrm{C}\subseteq\mathtt{AG}$, $\mathtt{LE}_{\mathrm{C}}=\mathtt{nei}^{\alpha}_{\mathrm{C}}$.
    A class of action models $\mathbf{AM}$ is $\alpha$-representable by a class of alpha neighborhood models $\mathbf{NM}^{\alpha}$ if (1) every action model in $\mathbf{AM}$ is $\alpha$-representable by an alpha neighborhood model in $\mathbf{NM}^{\alpha}$, and (2) every alpha neighborhood model in $\mathbf{NM}^{\alpha}$ $\alpha$-represents an action model in $\mathbf{AM}$.
@@
    For every class of action models $\mathbf{AM}$ and class of neighborhood models $\mathbf{NM}$, if $\mathbf{AM}$ is $z$-representable by $\mathbf{NM}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}$.
@@
    For every class of action models $\mathbf{AM}$ and class of alpha neighborhood models $\mathbf{NM}^{\alpha}$, if $\mathbf{AM}$ is $\alpha$-representable by $\mathbf{NM}^{\alpha}$, then for every formula $\phi$ in $\Phi$, $\phi$ is valid with respect to $\mathbf{AM}$ if and only if $\phi$ is valid with respect to $\mathbf{NM}^{\alpha}$.
Assume we have a coalition logic and the class of its models is $\mathbf{AM}$, a class of action models. By this result, if $\mathbf{AM}$ is representable by a class of neighborhood models $\mathbf{NM}$, then we can transform the action semantics of the coalition logic to the neighborhood semantics based on $\mathbf{NM}$.
From Fact 1, from every action model in $\mathtt{AM}$, we can get a neighborhood model $\mathtt{NM}$ representing $\mathtt{AM}$. Let $\mathbf{NM}$ be the collection of all these neighborhood models. Clearly, $\mathbf{AM}$ is representable by $\mathbf{NM}$. Note that this transformation is not very useful. To make a transformation meaningful, $\mathbf{NM}$ shall be defined directly.
In this paper, we mainly consider actual neighborhood models. In the sequel, by neighborhood models, we mean actual neighborhood models, unless explicitly stated.
3 Eight coalition logics determined by eight classes of
grand-coalition-first action models